09:00
R4 alpha background continues, currently R4_12
Temps OK, system checks unchanged from previous status (OK, aida16 clock and ADC issue)
10:30
R4 stopped and moved to the next run R5.
Lithium 3+ beam is not feasible at sufficient densities so we are switching to D+ at a beam energy of 0.75 MeV/u. A measurement of d+p at this energy would produce 3He at energies of ~900 keV.
The current threshold (slow comparator) has been 1 MeV (0x64). We lowered the thresholds in all FEE modules to 200 keV (0x14). ADC rates look fine in all FEEs (~tens of Hz) except in aida 15 and aida 16 which was ~several hundred kHz. Data rate to disk was 1.2 Mb/s
We increased the threshold in aida 15 and 16 only to 500 keV (0x32) and rates in aida 15 and 16 have dropped to <20 kHz. Data rate to disk now 220 kB/s.
11:00
R5 stopped. R6 started - background with reduced thresholds.
Attachment 1 - ADC data items statistics
Attachment 2 - Histogram browser statistics
Attachment 3 - Tape server
Attachment 4 - New Merger
13:12
R6 continues R6_9
Edits to the CARME sort code have been made. EyEx energy cut for online monitoring reduced to 100 keV. Multiplicity cut applies to lowEnergyEy, lowEnergyEx and lowEnergyXY histograms. Multiplicity cut considers only single strip events.
15:00
D+ beam is now in the ring and is being optimised by the beam operators. The CARME detectors are currently fullt out, but we can see some effects of the beam on the online monitoring code. |